# Barrier relaxations of the classical and quantum optimal transport problems

### v1

@@ 1.1. Discrete classical optimal transport problem
The discrete classical bi-partite optimal transport problem (BPOTP) was introduced by Hitchcock and Kantorovich [20, 21]. It is the following linear programming problem (LP). Let $\Delta^{n}\subset\mathbb{R}^{n}$ the simplex of column probablity vectors, and $[n]=\{1,\ldots,n\}$. For $\mathbf{x}=(x_{1},\ldots,x_{n})^{\top}\in\Delta^{n}$ denote by $E_{S}(\mathbf{x})=-\sum_{i=1}^{n}x_{i}\log x_{i}$ the Shannon entropy of $\mathbf{x}$. Assume that $\mathbf{p}_{i}\in\Delta^{n_{i}}$ for $i=1,2$. Let $P=(\mathbf{p}_{1},\mathbf{p}_{2})$ and $\mathrm{V}(P)$ be the convex set of all (coupling) probability matrices with marginals $\mathbf{p}_{1},\mathbf{p}_{2}$:
@@
(See part (a) of Lemma 2.1.) As $-\mathrm{E}_{S}$ is strictly convex, the minimal solution to (1.3) is unique. For a matrix $A=[a_{j_{1},j_{2}}]\in\mathbb{R}^{n_{1}\times n_{2}}$ denote by $\exp(\circ A)\in\mathbb{R}_{++}^{n_{1}\times n_{2}}$ the entry-wise exponential matrix $[\exp(a_{j_{1},j_{2}})]$. Then the minimal solution to the entropic relaxation of OTP is given by
@@ 1.3. Our contribution
(1.12) $(R^{n_{1}}\times\mathbb{R}^{n_{2}})_{++,b}=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}_{++}^{n_{1}}\times\mathbb{R}^{n_{2}}_{++},\mathbf{1}_{n_{1}}^{\top}\mathbf{x}=\mathbf{1}_{n_{2}}^{\top}\}.$
@@
$\sup_{\mathbf{x},\mathbf{y}\in\hat{\mathrm{D}}}\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}=\sup_{\mathbf{x},\mathbf{y}\in\mathrm{D}}\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}=$
One can apply the IPM method to the minimum problem (1.3) and minus the maximal problem (1.18). For $n_{1}=n_{2}=n$ the minimum problem (1.3) has $(n-1)^{2}$ parametrs, while the maximum problem has $12n-$ 1 parameters. Thus, the dual maximum characterization is superior.
@@
(1.21) $\kappa(C,R)=\sup_{U_{1}\in\mathrm{H}_{n_{1}},U_{2}\in\mathrm{H}_{n_{2}},C-U_{1}\otimes I_{n_{2}}-I_{n_{1}}\otimes U_{2}\succeq 0}\operatorname{Tr}(\rho^{1}U_{1})+\operatorname{Tr}(\rho^{2}U_{2}).$
@@
It is shown in [9, Theorem 3.2] that if $\rho^{1},\rho^{2}$ are positive definite then the supremum in (1.21) is achieved. However, if at least one of the density matrices is singular the supremum may not be achieved [9, Appendix B.3].
@@ 2.1. An entropic relaxation of the classical bi-partite optimal transport problem
Let $\mathbf{p}_{i}\in\Delta^{n_{i}},n_{i}>1,i\in[2]$ be positive probability vectors, and assume that $\mathrm{U}(P)$ be defined by (1.1).
@@
    The unique minimizing matrix $U\in\mathrm{U}_{o}(P)$ for (1.3) is of the form (1.5), where $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ are unique.
@@
(b) The quantity $\tau(C,P,\varepsilon)$ is a constrained minimum problem. As $-E_{S}$ is strictly convex, and a directional derivative of $-E_{S}$ toward the entrior of $\mathrm{U}(P)$ at any boundary point is $\infty$- ∞ it follows that the unique minimum point $U=[u_{j_{1},j_{2}}]$ is in $\mathrm{V}_{o}(P)$. That is, $U\in\mathbb{R}_{++}^{n_{1}\times n_{2}}$. Recall that $\mathrm{V}(P)$ is given by $n_{1}+n_{2}$ equalities given by (1.1). (Note we have the equality that the sum of the rows are equal to the sum of the columns of $V\in\mathbb{R}^{n_{1}\times n_{2}}$.) Hence, we can use the Lagrange multiplies: $\lambda=(\lambda_{1},\ldots,\lambda_{n_{1}})^{\top},\nu=(\nu_{1},\ldots,\nu_{n_{2}})^{\top}$ corresponding to the row and column sums of $V\in\mathrm{V}(P)$. Apply the Lagrange multipliers to deduce [14]
@@
Observe that if we replace $\lambda_{j_{1}}$ and $\mu_{j_{2}}$ by $\lambda_{j_{1}}+t$ and $\mu_{j_{2}}-t$ for $j_{1}\in[n_{1}],j_{2}\in[n_{2}]$ we still have the above equalites. Thus, we can choose a unique $t$ such that $\sum_{j_{1}=1}^{n_{1}}(\lambda_{j_{1}}-\varepsilon+t)=\sum_{j_{2}=1}^{n_{2}}(\mu_{j_{2}}-t)$. This shows that the unique minimal $U$ is of the form (1.5). It is left to show that $\mathbf{u}_{1},\mathbf{u}_{2}$ are unique. Since the minimal matrix $U=[u_{j_{1},j_{2}}]$ is unique, and $C$ is known, we obtain that $\mathbf{u}_{1},\mathbf{u}_{2}$ satisfy the equalities:
(2.1) $u_{j_{1},1}+u_{j_{2},2}=a_{j_{1},j_{2}},\quad j_{1}\in[n_{1}],j_{2}\in[n_{2}],$
where $A=[a_{j_{1},j_{2}}]$ is known. Recall that $\mathbf{1}_{n_{1}}^{\top}\mathbf{u}_{1}=1_{n_{2}}^{\top}\mathbf{u}_{2}=t$. Sum up the $n_{1}n_{2}$ equalities in (2.1) to deduce that $(n_{1}+n_{2})t=\mathbf{1}_{n_{1}}^{\top}A\mathbf{1}_{n_{2}}$. Sum up the equalites in (2.1) on either $j_{2}$ or $j_{1}$ to obtain
$\mathbf{u}_{1}=\frac{1}{n_{2}}(A\mathbf{1}_{n_{2}}-t\mathbf{1}_{n_{1}}),\mathbf{u}_{2}=\frac{1}{n_{1}}(A^{\top}\mathbf{1}_{n_{1}}-t\mathbf{1}_{n_{2}}),t=\frac{\mathbf{1}_{n_{1}}^{\top}A\mathbf{1}_{n_{2}}}{n_{1}+n_{2}}.$
@@
The problem with entropic relaxation is that we need to rescale the matrix $\exp(\circ(-\frac{C}{\varepsilon}))$. Then, if $\frac{\min_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|}{\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|}>0$ but small, and $\varepsilon$ is small ($=10^{-8})$, then many entries of $\exp(\circ(-\frac{C}{\varepsilon}))$ would be treated as zero entries in the scaling algorithm.
@@ 2.2. A barrier relaxation of the classical bi-partite optimal transport problem
Let $(\mathbf{p}_{1},\mathbf{p}_{2})\in(\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}})_{++,b}$. Assume that $\mathrm{V}(P)$ be defined by (1.1).
@@
Hence, $(\mathbf{u}_{1},\mathbf{u}_{2})$ is a maximal point in $\hat{\mathrm{D}}$. As $\mathbf{1}_{n_{1}}^{\top}\mathbf{u}_{1}=\mathbf{1}_{n_{2}}^{\top}\mathbf{u}_{2}$, it follows that $(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathrm{D}$. As $\varphi$ is strictly concave in $\mathrm{D}$, it follows that $(\mathbf{u}_{1},\mathbf{u}_{2})$ is the unique maximal point in $\mathrm{D}$. Let $(\mathbf{x}^{\star},\mathbf{y}^{\star})\in\hat{\mathrm{D}}\setminus\mathrm{D}$ be a maximum point in $\hat{\mathrm{D}}$. Then $(\mathbf{x}^{\star},\mathbf{y}^{\star})+t(\mathbf{1}_{n_{1}},-\mathbf{1}_{n_{2}})$ is also a maximal point in $\hat{\mathrm{D}}$. There exists a unique $t^{\star}$ such that $(\mathbf{x}^{\star},\mathbf{y}^{\star})+t^{\star}(\mathbf{1}_{n_{1}},-\mathbf{1}_{n_{2}})\in\mathrm{D}$. Hence, $(\mathbf{x}^{\star},\mathbf{y}^{\star})+t^{\star}(\mathbf{1}_{n_{1}},-\mathbf{1}_{n_{2}})=(\mathbf{u}_{1},\mathbf{u}_{2})$. Therefore, the maximal point of $\varphi$ in $\hat{\mathrm{D}}$.
@@ 2.3. The interior point method for finding a barrier approximation of $\tau(C,P)$
(2.4) $\frac{n_{1}\cdots n_{d}}{\max_{i\in[d]}n_{i}}\sum_{i=1}^{d}\|\mathbf{x}_{i}\|^{2}\leq\sum_{j_{1}=\cdots=j_{d}=1}^{n_{1},\cdots,n_{d}}(x_{j_{1},1}+\cdots+x_{j_{d},d})^{2}\leq$
$\sum_{i=1}^{d}\frac{dn_{1}\cdots n_{d}}{n_{i}}\|\mathbf{x}_{i}\|^{2}\leq\frac{dn_{1}\cdots n_{d}}{\max_{i\in[d]}n_{i}}\sum_{i=1}^{d}\|\mathbf{x}_{i}\|^{2}.$
@@
$\sum_{i=1}^{d}\frac{n_{1}\cdots n_{d}}{n_{i}}\|\mathbf{x}_{i}\|^{2}+2\sum_{1\leq p<q\leq d}(\mathbf{1}_{n_{p}}^{\top}\mathbf{x}_{n_{p}})(\mathbf{1}_{n_{q}}^{\top}\mathbf{x}_{n_{q}})$
@@
Let $2\leq n_{1},n_{2}\in\mathbb{N}$, and $0\neq C=[c_{j_{1},j_{2}}]\in\mathbb{R}^{n_{1}\times n_{2}}$. Denote
@@
    Let $\mathbf{p}_{1},\mathbf{p}_{2}$ are positive probability vectors. Assume that $C(\mathbf{x},\mathbf{y})$ is defined in (1.1), and suppose that
@@
    (2.9) $\sum_{j_{1}=j_{2}=1}^{n_{1},n_{2}}(x_{j_{1}}+y_{j_{2}})^{2}<n_{1}n_{2}\left(\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\right)^{2},$
@@
    The interior of the ball $\mathrm{B}((\mathbf{x}_{0},\mathbf{y}_{0}),r)=\{\|\mathbf{x}-\mathbf{x}_{0}\|^{2}+\|\mathbf{y}-\mathbf{y}_{0}\|^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((\mathbf{x}_{0},\mathbf{y}_{0}),r)$, contains $(\mathbf{x},\mathbf{y})$ that satisfies the conditoins (2.8).
@@
Consider the minimum problem (1.2). Then $\mathbf{b}=(\mathbf{p}_{1},\mathbf{p}_{2}),\mathbf{w}=(\mathbf{x},\mathbf{y})$ viewed as column vectors in $\mathbb{R}^{n_{1}+n_{2}}$. Thus, $\mathbf{b}^{\top}\mathbf{w}=\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}$. We claim that the condition $\mathbf{c}-A^{\top}\mathbf{w}\geq 0$ is the condition $C-\mathbf{x}\mathbf{1}_{n_{2}}^{\top}-\mathbf{1}_{n_{1}}\mathbf{y}^{\top}\geq 0$. To show that, observe the following implications
@@
Hence , the first equality line of (1.20) holds. Therefore, the second equality of (1.20) holds. Without loss of generality we can replace $(\mathbf{x},\mathbf{y})$ by $(\mathbf{x}+t\mathbf{1}_{n_{1}},\mathbf{y}-t\mathbf{1}_{n_{2}})$. Hence, the third equality of (1.20) holds. Furthermore, we can assume that there is a maximum solution $(\mathbf{u}_{1},\mathbf{u}_{2})$ to (1.20) such that $\mathbf{1}_{n_{1}}^{\top}\mathbf{u}_{1}=1_{n_{2}}^{\top}\mathbf{u}_{2}$.
@@
$\frac{c_{\min}-\operatorname{Tr}C^{\top}(\mathbf{p}_{1}\mathbf{p}_{2}^{\top})+p_{1,k}p_{2,l}c_{p,q}}{p_{1,k}p_{2,l}}\leq x_{p}+y_{q}\,\forall(p,q)\in[n_{1}]\times[n_{2}]\Rightarrow$
@@
$\|\mathbf{x}-\mathbf{x}_{0}\|^{2}+\|\mathbf{y}-\mathbf{y}_{0}\|^{2}\leq(\|\mathbf{x}\|+\|\mathbf{x}_{0}\|)^{2}+(\|\mathbf{x}\|+\|\mathbf{x}_{0}\|)^{2}<$
$2(\|\mathbf{x}\|^{2}+\|\mathbf{x}_{0}\|^{2}+\|\mathbf{y}\|^{2}+\|\mathbf{y}_{0}\|^{2})<\frac{n_{1}n_{2}}{n_{1}+n_{2}}(-1+c_{\min})^{2}+$
@@
Let $\mathbf{p}_{i}\in\Delta^{n_{i}},i\in[2]$ be positive probability vectors. Consider the minimum (1.2), which corresponds to the maximum dual problem (1.20). Apply the IPM short step interior path algorithm with the barrier (2.12) to the $-\min$ problem as in (2.2), where $r$ is given by (3.18). Let $c=\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|$. Choose a starting point given by (3.17). The CPA algorithm finds the value $\tau(C,P)$ within precision $\delta>0$ in
@@
Here, $\textrm{sym}((\mathbf{x}_{0},\mathbf{y}_{0}),\mathrm{D}_{1})$ is bounded below the ration of the minimum distance and the maximum distance of $(\mathbf{x}_{0},\mathbf{y}_{0})$ to the boundary. Part (e) of Lemma 2.4 yields that $\textrm{sym}((\mathbf{x}_{0},\mathbf{y}_{0}),\mathrm{D}_{1})\geq\sqrt{2}r$. This shows that the number of iterations to find the value $\tau(C,P)$ within precision $\delta>0$ is given by (2.15).

### v2

@@ 1.1. Discrete classical optimal transport problem
The discrete classical bi-partite optimal transport problem (BPOTP) was introduced by Hitchcock and Kantorovich [20, 21]. It is the following linear programming problem (LP). Let $\Delta^{n}\subset\mathbb{R}^{n}$ the simplex of column probablity vectors, and $n]=\{1,\ldots,n\}$. For $\mathbf{x}=(x_{1},\ldots,x_{n})^{\top}\in\Delta^{n}$ denote by $E_{S}(\mathbf{x})=-\sum_{i=1}^{n}x_{i}\log x_{i}$ the Shannon entropy of $\mathbf{x}$. Assume that $\mathbf{p}_{i}\in\Delta^{n_{i}}$ for $i=1,2$. Let $P=(\mathbf{p}_{1},\mathbf{p}_{2})$ and $\mathrm{V}(P)$ be the convex set of all (coupling) probability matrices with marginals $\mathbf{p}_{1},\mathbf{p}_{2}$:
@@
(See part (a) of Lemma 2.1.) As $\mathrm{E}_{S}$ is strictly convex, the minimal solution to (1.3) is unique. For a matrix $A=[a_{j_{1},j_{2}}]\in\mathbb{R}^{n_{1}\times n_{2}}$ denote by $\exp(\circ A)\in\mathbb{R}_{++}^{n_{1}\times n_{2}}$ the entry-wise exponential matrix $\exp(a_{j_{1},j_{2}})]$. Then the minimal solution to the entropic relaxation of OTP is given by
@@ 1.3. Our contribution
(1.12) $(R^{n_{1}}\times\mathbb{R}^{n_{2}})_{++,b}=\{(\mathbf{x},\mathbf{y})\in\mathbb{R}_{++}^{n_{1}}\times\mathbb{R}^{n_{2}}_{++},\mathbf{1}_{n_{1}}^{\top}\mathbf{x}=\mathbf{1}_{n_{2}}^{\top}\mathbf{y}\}.$
@@
$\sup_{\mathbf{x},\mathbf{y}\in\hat{\mathrm{D}}}\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}=\sup_{\mathbf{x},\mathbf{y}\in\mathrm{D}}\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}$
One can apply the IPM method to the minimum problem (1.3) and minus the maximal problem (1.18). For $n_{1}=n_{2}=n$ the minimum problem (1.3) has $(n-1)^{2}$ parametrs, while the maximum problem has 2n12n-1 parameters. Thus, the dual maximum characterization is superior.
@@
(1.21) $\kappa(C,R)=\sup_{U_{1}\in\mathrm{H}_{n_{1}},U_{2}\in\mathrm{H}_{n_{2}},C-U_{1}\otimes I_{n_{2}}-I_{n_{1}}\otimes U_{2}\succeq 0}\operatorname{Tr}(\rho_{1}U_{1})+\operatorname{Tr}(\rho_{2}U_{2}).$
@@
It is shown in [9, Theorem 3.2] that if $\rho_{1},\rho_{2}$ are positive definite then the supremum in (1.21) is achieved. However, if at least one of the density matrices is singular the supremum may not be achieved [9, Appendix B.3].
@@ 2.1. An entropic relaxation of the classical bi-partite optimal transport problem
Let $\mathbf{p}_{i}\in\Delta^{n_{i}},n_{i}>1,i\in[2]$ be positive probability vectors, and assume that $\mathrm{V}(P)$ be defined by (1.1).
@@
    The unique minimizing matrix $U\in\mathrm{V}_{o}(P)$ for (1.3) is of the form (1.5), where $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ are unique.
@@
(b) The quantity $\tau(C,P,\varepsilon)$ is a constrained minimum problem. As $E_{S}$ is strictly convex, and a directional derivative of $E_{S}$ toward the entrior of $\mathrm{U}(P)$ at any boundary point is -$\infty$ it follows that the unique minimum point $U=[u_{j_{1},j_{2}}]$ is in $\mathrm{V}_{o}(P)$. That is, $U\in\mathbb{R}_{++}^{n_{1}\times n_{2}}$. Recall that $\mathrm{V}(P)$ is given by $n_{1}+n_{2}$ equalities given by (1.1). (Note we have the equality that the sum of the rows are equal to the sum of the columns of $V\in\mathbb{R}^{n_{1}\times n_{2}}$.) Hence, we can use the Lagrange multiplies: $\lambda=(\lambda_{1},\ldots,\lambda_{n_{1}})^{\top},\nu=(\nu_{1},\ldots,\nu_{n_{2}})^{\top}$ corresponding to the row and column sums of $V\in\mathrm{V}(P)$. Apply the Lagrange multipliers to deduce [14]
@@
Observe that if we replace $\lambda_{j_{1}}$ and $\mu_{j_{2}}$ by $\lambda_{j_{1}}+t$ and $\mu_{j_{2}}-t$ for $j_{1}\in[n_{1}],j_{2}\in[n_{2}]$ we still have the above equalites. Thus, we can choose a unique $t$ such that $\sum_{j_{1}=1}^{n_{1}}(\lambda_{j_{1}}-\varepsilon+t)=\sum_{j_{2}=1}^{n_{2}}(\mu_{j_{2}}-t)$, where $u_{j_{1},1}=(\lambda_{j_{1}}-\varepsilon+t)/\varepsilon,u_{j_{2},2}=(\mu_{j_{2}}-t)/\varepsilon$. This shows that the unique minimal $U$ is of the form (1.5). It is left to show that $\mathbf{u}_{1},\mathbf{u}_{2}$ are unique. Since the minimal matrix $U=[u_{j_{1},j_{2}}]$ is unique, and $C$ is known, we obtain that $\mathbf{u}_{1},\mathbf{u}_{2}$ satisfy the equalities:
(2.1) $u_{j_{1},1}+u_{j_{2},2}=\log u_{j_{1},j_{2}}+c_{j_{1},j_{2}}/\varepsilon=a_{j_{1},j_{2}},\quad j_{1}\in[n_{1}],j_{2}\in[n_{2}],$
where $A=[a_{j_{1},j_{2}}]$ is known. Recall that $\mathbf{1}_{n_{1}}^{\top}\mathbf{u}_{1}=1_{n_{2}}^{\top}\mathbf{u}_{2}=s$. Sum up the $n_{1}n_{2}$ equalities in (2.1) to deduce that $(n_{1}+n_{2})s=\mathbf{1}_{n_{1}}^{\top}A\mathbf{1}_{n_{2}}$. Sum up the equalites in (2.1) on either $j_{2}$ or $j_{1}$ to obtain
$\mathbf{u}_{1}=\frac{1}{n_{2}}(A\mathbf{1}_{n_{2}}-s\mathbf{1}_{n_{1}}),\quad\mathbf{u}_{2}=\frac{1}{n_{1}}(A^{\top}\mathbf{1}_{n_{1}}-s\mathbf{1}_{n_{2}}),\quad s=\frac{\mathbf{1}_{n_{1}}^{\top}A\mathbf{1}_{n_{2}}}{n_{1}+n_{2}}.$
@@
The problem with entropic relaxation is that we need to rescale the matrix $\exp(\circ(-\frac{C}{\varepsilon}))$. Then, if $\frac{\min_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|}{\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|}>0$ but small, and $\varepsilon$ is small ($10^{-8})$, then many entries of $\exp(\circ(-\frac{C}{\varepsilon}))$ would be treated as zero entries in the scaling algorithm.
@@ 2.2. A barrier relaxation of the classical bi-partite optimal transport problem
Let $\mathbf{p}_{1}\in\mathbb{R}^{n_{1}},\mathbf{p}_{2}\in\mathbb{R}^{n_{2}}$ be positive probability vectors. Assume that $\mathrm{V}(P)$ be defined by (1.1).
@@
Hence, $(\mathbf{u}_{1},\mathbf{u}_{2})$ is a maximal point in $\hat{\mathrm{D}}$. As $\mathbf{1}_{n_{1}}^{\top}\mathbf{u}_{1}=\mathbf{1}_{n_{2}}^{\top}\mathbf{u}_{2}$, it follows that $(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathrm{D}$. As $\varphi$ is strictly concave in $\mathrm{D}$, it follows that $(\mathbf{u}_{1},\mathbf{u}_{2})$ is the unique maximal point in $\mathrm{D}$. Let $(\mathbf{x}^{\star},\mathbf{y}^{\star})\in\hat{\mathrm{D}}\setminus\mathrm{D}$ be a maximum point in $\hat{\mathrm{D}}$. Then $(\mathbf{x}^{\star},\mathbf{y}^{\star})+t(\mathbf{1}_{n_{1}},-\mathbf{1}_{n_{2}})$ is also a maximal point in $\hat{\mathrm{D}}$. There exists a unique $t^{\star}$ such that $(\mathbf{x}^{\star},\mathbf{y}^{\star})+t^{\star}(\mathbf{1}_{n_{1}},-\mathbf{1}_{n_{2}})\in\mathrm{D}$. Hence, $(\mathbf{x}^{\star},\mathbf{y}^{\star})+t^{\star}(\mathbf{1}_{n_{1}},-\mathbf{1}_{n_{2}})=(\mathbf{u}_{1},\mathbf{u}_{2})$. Therefore, maximal points of $\varphi$ in $\hat{\mathrm{D}}$ are given by (1.19).
@@ 2.3. The interior point method for finding a barrier approximation of $\tau(C,P)$
(2.4) $\sum_{i=1}^{d}\|\mathbf{x}_{i}\|^{2}\leq\frac{\max_{i\in[d]}n_{i}}{n_{1}\cdots n_{d}}\sum_{j_{1}=\cdots=j_{d}=1}^{n_{1},\cdots,n_{d}}(x_{j_{1},1}+\cdots+x_{j_{d},d})^{2}\leq$
$(\max_{i\in[d]}n_{i})\sum_{i=1}^{d}\frac{d}{n_{i}}\|\mathbf{x}_{i}\|^{2}$
@@
$\sum_{i=1}^{d}\frac{n_{1}\cdots n_{d}}{n_{i}}\|\mathbf{x}_{i}\|^{2}+2\sum_{1\leq p<q\leq d}\frac{n_{1}\cdots n_{d}}{n_{p}n_{q}}(\mathbf{1}_{n_{p}}^{\top}\mathbf{x}_{n_{p}})(\mathbf{1}_{n_{q}}^{\top}\mathbf{x}_{n_{q}})$
@@
Let $2\leq n_{1},n_{2}\in\mathbb{N}$, $0\neq C=[c_{j_{1},j_{2}}]\in\mathbb{R}^{n_{1}\times n_{2}}$, and $D$ be given by (1.15). Denote
@@
    Let $\mathbf{p}_{1},\mathbf{p}_{2}$ be positive probability vectors. Assume that $C(\mathbf{x},\mathbf{y})$ is defined in (1.15), and suppose that
@@
    (2.9) $0\leq c_{j_{1},j_{2}}-x_{j_{1}}-y_{j_{2}}<\frac{c_{\max}-c_{\min}}{p_{j_{1},1}p_{j_{2},2}}\leq$
    $\frac{2c}{p_{j_{1},1}p_{j_{2},2}}\leq\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})},$
    $\sum_{j_{1}=j_{2}=1}^{n_{1},n_{2}}(x_{j_{1}}+y_{j_{2}})^{2}<n_{1}n_{2}\left(\frac{2c}{(\min_{j_{1}\in[n_{1}]}p_{j_{1},1})(\min_{j_{2}\in[n_{2}]}p_{j_{2},2})}\right)^{2},$
@@
    The interior of the ball $\mathrm{B}((\mathbf{x}_{0},\mathbf{y}_{0}),r)=\{\|\mathbf{x}-\mathbf{x}_{0}\|^{2}+\|\mathbf{y}-\mathbf{y}_{0}\|^{2}\leq r^{2}\},$ denoted as $\mathrm{B}_{o}((\mathbf{x}_{0},\mathbf{y}_{0}),r)$, contains all $(\mathbf{x},\mathbf{y})$ that satisfies the conditoins (2.8).
@@
Consider the minimum problem (1.2). Then $\mathbf{b}=(\mathbf{p}_{1},\mathbf{p}_{2}),\mathbf{w}=(\mathbf{x},\mathbf{y})$ viewed as column vectors in $\mathbb{R}^{n_{1}+n_{2}}$. Thus, $\mathbf{b}^{\top}\mathbf{w}=\mathbf{p}_{1}^{\top}\mathbf{x}+\mathbf{p}_{2}^{\top}\mathbf{y}$. Observe first that the condition $\mathbf{c}-A^{\top}\mathbf{w}\geq 0$ is the condition $C-\mathbf{x}\mathbf{1}_{n_{2}}^{\top}-\mathbf{1}_{n_{1}}\mathbf{y}^{\top}\geq 0$. To show that, observe the following implications
@@
Furthermore, the duality yields first equality line of (1.20). Therefore, the second equality of (1.20) holds. Without loss of generality we can replace $(\mathbf{x},\mathbf{y})$ by $(\mathbf{x}+t\mathbf{1}_{n_{1}},\mathbf{y}-t\mathbf{1}_{n_{2}})$. Hence, the third equality of (1.20) holds. Furthermore, we can assume that there is a maximum solution $(\mathbf{u}_{1},\mathbf{u}_{2})$ to (1.20) such that $\mathbf{1}_{n_{1}}^{\top}\mathbf{u}_{1}=1_{n_{2}}^{\top}\mathbf{u}_{2}$.
@@
$0\leq c_{k,l}-x_{p}-y_{q}\leq\frac{\operatorname{Tr}C^{\top}(\mathbf{p}_{1}\mathbf{p}_{2}^{\top})-c_{\min}}{p_{1,k}p_{2,l}}\leq\frac{c_{\max}-c_{min}}{p_{1,k}p_{2,l}}\leq\frac{2c}{p_{1,k}p_{2,l}}.$
The last inequality yields the first inequalities in (2.9). Next observe that the above inequalities yield:
$-\frac{2c}{p_{1,k}p_{2,l}}<\frac{c_{\min}-\operatorname{Tr}C^{\top}(\mathbf{p}_{1}\mathbf{p}_{2}^{\top})+p_{1,k}p_{2,l}c_{k,l}}{p_{1,k}p_{2,l}}\leq x_{k}+y_{l}\leq c_{k,l}\Rightarrow$
@@
$\|\mathbf{x}-\mathbf{x}_{0}\|^{2}+\|\mathbf{y}-\mathbf{y}_{0}\|^{2}\leq(\|\mathbf{x}\|+\|\mathbf{x}_{0}\|)^{2}+(\|\mathbf{y}\|+\|\mathbf{y}_{0}\|)^{2}<$
$2(\|\mathbf{x}\|^{2}+\|\mathbf{x}_{0}\|^{2}+\|\mathbf{y}\|^{2}+\|\mathbf{y}_{0}\|^{2})<\frac{2n_{1}n_{2}}{n_{1}+n_{2}}(-1+c_{\min})^{2}+$
@@
Let $\mathbf{p}_{i}\in\Delta^{n_{i}},i\in[2]$ be positive probability vectors. Consider the minimum (1.2), which corresponds to the maximum dual problem (1.20). Apply the IPM short step interior path algorithm with the barrier (2.12) to the $\min$ problem as in (2.2), where $r$ is given by (2.11). Let $c=\max_{j_{1}\in[n_{1}],j_{2}\in[n_{2}]}|c_{j_{1},j_{2}}|$. Choose a starting point $(\mathbf{x}_{0},\mathbf{y}_{0})$ given by (2.10). The CPA algorithm finds the value $\tau(C,P)$ within precision $\delta>0$ in
@@
Here, $\textrm{sym}((\mathbf{x}_{0},\mathbf{y}_{0}),\mathrm{D}_{1})$ is bounded below by the ratio of the minimum distance and the maximum distance of $(\mathbf{x}_{0},\mathbf{y}_{0})$ to the boundary. Part (e) of Lemma 2.4 yields that $\textrm{sym}((\mathbf{x}_{0},\mathbf{y}_{0}),\mathrm{D}_{1})\geq\sqrt{2}r$. This shows that the number of iterations to find the value $\tau(C,P)$ within precision $\delta>0$ is given by (2.15).